HETAV PATEL

Grade 11,

An object is thrown vertically up from the ground passes the height 5 m twice in an interval of 10 secs. what is the time of flight? 1)28 0.5 2)86 0.5 3)104 0.5 4)72 0.5

An object is thrown vertically up from the ground passes the height 5 m twice in an interval of 10 secs. what is the time of flight?
1)280.5
2)860.5
3)1040.5
4)720.5

Grade:11

2 Answers

bharat bajaj IIT Delhi
9 years ago
s = ut + 1/2. a.t^2 .....
5 = ut - 5t^2 .....
5t^2 - ut + 5 = 0 ...
The roots of this equation gives us the time at which particle crosses the height 5 m....If we calculate the difference of roots, it tells us the interval which we know is 10 s...
Say roots are a and b....a-b = ( (a+b)^2 - 4ab ) ^0.5....a - b = ( u^2/25 - 4)^0.5....100 = u^2/25 - 4..
u = 5.(104)^0.5...Total time of flight = 2u/a = (104)^0.5..THanks
Bharat BajajIIT Delhi
askiitians faculty
Arun Kumar IIT Delhi
9 years ago
Hello
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty